Economic Help

The price elasticity of demand is inelastic for gasoline and elastic for tablets. Suppose that technological advance doubles the supply of both products (that is, the quantity supplied at each price is twice what it was).

a. What happens to the equilibrium price and quantity in each market? Use a supply-and-demand graph for both gasoline and tablets and analyze which product experiences a larger change in price and which product experiences a larger change in quantity.

b. What happens to total revenue for each product? Briefly explain.
Answer #1

A.

1. When the price elasticity is inelastic for gasoline supply.

Gasoline product will experience large change in price and small change in quantity as demand is inelastic for the gasoline.

2.when the price elasticity is elastic for tablets.

Here the price for tablets will smaller change and quantity of the products will change higher as it is elastic demand for the tablets.

B.

Revenue for gasoline.

Revenue = P*Q

So here before change in supply total revenue was=

=Price * quantity

=15*100

=1500

Revenue after supply change

=5*200

=1000

So if the demand for the gasoline product is inelastic and if the supply changes the revenue will decrease.

Revenue for tablets.

Revenue before change in supply

=15*100

=1500.

Revenue after change in supply

=12*200

=2400

So here as demand is elastic for tablets if we increases the supply, revenue will be also increased.